In a girl guide camp 80girls had food for 15 days, 40 more jioned them after 3 days, for how mnay days the remaining food will suffice.
: In a girl guide camp 80girls had food for 15 days, 40 more jioned them after 3 days, for how mnay days the remaining food will suffice.

Answer by checkley77(3639)

80*15=1200 number of girl days total.
80*3=240 girl days used in the first 3 days.
1200-240=960 girl days left.
80+40=120 girls for the remainder of the camp.
960/120=8 more days all 120 girls can be fed.